One alloy is 70% lead and 30% zinc, and another is 40% lead and 60% zinc. Find the amount of each alloy needed to make 240kg of an alloy that is 50% lead and 50% zinc. I am having trouble finding the equation to this problem.
Answer by xinxin(76)

Suppose the amount of the first alloy needed is X kg, the second is Y kg:
0.7X + 0.4Y = 240 * 0.5 = 120
0.3X + 0.6Y = 240 * 0.5 =120
X = 80 , Y = 160